10) Pisa Pizza, a seller of frozen pizza, is considering introducing a healthier version of its...

10)

Pisa Pizza, a seller of frozen pizza, is considering introducing a healthier version of its pizza that will be low in cholesterol and contain no trans fats. The firm expects that sales of the new pizza will be $21 million per year. While many of these sales will be to new customers, Pisa Pizza estimates that 45% will come from customers who switch to the new, healthier pizza instead of buying the original version.?? a. Assume customers will spend the same amount on either version. What level of incremental sales is associated with introducing the new pizza? b. Suppose that 44% of the customers who will switch from Pisa Pizza's original pizza to its healthier pizza will switch to another brand if Pisa Pizza does not introduce a healthier pizza. What level of incremental sales is associated with introducing the new pizza in this case? a. Assume customers will spend the same amount on either version. What level of incremental sales is associated with introducing the new pizza? The incremental sales are $ -------- million. (Round to two decimal places.)

12)

A bicycle manufacturer currently produces 237,000 units a year and expects output levels to remain steady in the future. It buys chains from an outside supplier at a price of $1.90 a chain. The plant manager believes that it would be cheaper to make these chains rather than buy them. Direct? in-house production costs are estimated to be only $1.50 per chain. The necessary machinery would cost $273,000 and would be obsolete after ten years. This investment could be depreciated to zero for tax purposes using a? ten-year straight-line depreciation schedule. The plant manager estimates that the operation would require $53,000 of inventory and other working capital upfront? (year 0), but argues that this sum can be ignored since it is recoverable at the end of the ten years. Expected proceeds from scrapping the machinery after ten years are $20,475. If the company pays tax at a rate of 35% and the opportunity cost of capital is 15%?, what is the net present value of the decision to produce the chains? in-house instead of purchasing them from the? supplier? Project the annual free cash flows (FCF?) of buying the chains.

The annual free cash flows for years 1 to 10 of buying the chains is $ -------.

? (Round to the nearest dollar. Enter a free cash outflow as a negative? number.)
Answer #1

Question -1]

The Total Sales after introduction of new pizza = $ 21 Million

Sales due to new customer = 55% of $ 21 Million

a) Incremental sales due to introduction of New Pizza = $ 11.55 Million

b) Had 44% existing customers switched due to non introduction of new pizza, then in such case the incremental sales = (55 % + 44% ) i.e. 99% of $ 21 Million

Incremental sales = $ 20.79 Million

Question -2]

NPV is calculated using following formula

NPV = - Capital Investment + Cash inflow for first year /(1 +R) + cash inflow for second year / (1+R)^2 + So on .......................Till completion of 10th year.

Given Data –

Rate = 15%

Total Investment required = 273000

Number of units produced = 237000

Total Annual Profit = Number of units produced x (old cost – new cost)

= 237000 x (1.90 – 1.50)

= 94800

Depreciation = (273000 – 20475) / 10 = 25252.5 (annual)

Salvage value = 20475

Now,

Annual Free Cash Flow from 1st to 9th year = Annual profit x (1 – Tax percent) + Depreciation

= 94800 x (1 – 0.35) + 25252.5 = 86872.5

And,

Annual Free Cash Flow for 10th year = Annual profit x (1 – Tax percent) + Depreciation + Salvage Value

= 94800 x (1 – 0.35) + 25252.5 + 20475 = 107347.5

Substitute all values in NPV equation given above

NPV = -273000 + 86872.5 / (1.15) + ......................................+ 107347.5 /(1.15)^10

NPV = 168054.08 (In house production advised)